### v1

# Representative electricity price profiles for European day-ahead and intraday spot markets

Abstract
We propose a method to construct representative price profiles of the day-ahead (DA) and the intraday (ID) electricity spot markets and use this method to provide examples of ready-to-use price data sets. In contrast to common scenario generation approaches, the method is deterministic and relies on a small number of degrees of freedom, with the aim to be well-defined and easy to use. We thereby target an enhanced comparability of future research studies on demand-side management and energy cost optimization. We construct the price profiles based on historical time series from the spot markets of interest, e.g., European Power Exchange (EPEX) spot. To this end, we extract key price components from the data while also accounting for known dominant mechanisms in the price variation. Further, the method is able to preserve key statistical features of the historical data (e.g., mean and standard deviation) when constructing the benchmark profile. Finally, our approach ensures comparability of ID and DA price profiles by design, as their cumulative (integral) price can be made identical if needed.
Keywords: Pricing, Energy markets, Double peak profile, Demand response, Flexible operation
1   Introduction
Optimization of energy costs has long been acknowledged as a topic of interest for the economic design and operation of chemical processes and energy systems (Friedler,, 2010). Demand-side management (DSM) and demand response (DR) are energy-flexible operation paradigms that have have raised interest of industrial consumers to actively participate in the energy spot markets by adjustment of electricity demand (Pinson et al.,, 2014; Mitsos et al.,, 2018). With electricity prices on European DA spot markets having reached temporary peak values as high as 4 000 €/MWh (Nord Pool,, 2023), the importance of industrial DSM is expected to grow further in the near future. Thus, economic studies on planning and scheduling of industrial processes necessitate a systematic way to select electricity price scenarios. On the other hand, the process of selecting a price scenario should not distract from the main purpose of the paper. Herein, we thus propose a simple and intuitive approach to determine representative price profiles.
A common approach to select an electricity price profile is the use of a recent historical period, e.g., 24 hours, that exhibits a typical behavior in terms of the price range and transient pattern (e.g., Dalle Ave et al., (2019)). Especially when design and planning studies are performed, multiple historical price profiles can be considered to capture the effect of different fluctuation patterns (e.g., Fürsch et al., (2013)). Long horizons of historical data (e.g., up to one year) may be applied if needed (Meese et al.,, 2016). However, most commonly single-day or single-week price profiles are utilized. The selection of historical prices usually follows a non-systematic procedure, wherefore the representativity and generalizability of case study results from the literature is frequently obscure.
One way to improve the representative character of scenarios is to construct average price profiles over a historical period that respects fundamental characteristics of the spot market price profiles. Knittel & Roberts, (2005) identify electricity price trends on different time scales highlighting recurring patterns and characteristic values of the price time series. Rahimiyan & Baringo, (2016) perform a statistical analysis on historical DA and ID price data to quantify the price correlations. Schäfer et al., (2020) and Germscheid et al., (2022) utilize Fourier transforms to analyze principal frequencies of historical DA prices and ID-DA price deviations, respectively, identifying dominant modes at periods of 0.5 day and 1 day. By averaging one year of historical DA and ID price data, significant time-series frequencies, such as daily or weekly, are taken into account to construct a single profile. These periods correspond to significant price patterns and price fluctuations that can be exploited by most processes. This method, though straightforward, may over-smooth the data and potentially obscure certain less prominent trends.
A collection of more sophisticated approaches to determine representative periods in time-series data is presented in the review by Hoffmann et al., (2020). Clustering historical data provides a set of profiles that represents the entity of the historical data, e.g., by selecting the centroid of each cluster (Poncelet et al.,, 2017). Teichgraeber & Brandt, (2018, 2019) employ clustering to generate representative single-day profiles of one-year DA price data. Although the method reduces the number of scenarios, compared to considering the full collection of available data, multiple optimization problems need to be solved and the results are only optimal with respect to the chosen clustering criteria.
Finally, we mention probabilistic forecasting techniques, which exploit correlations of predicted parameters (e.g., weather forecasts) and are widely used to explain price behavior (Nowotarski & Weron,, 2018). Following Monteiro et al., (2018) and Weron, (2014), these methods can be categorized into prediction interval, density, and threshold forecasting. Univariate and multivariate probabilistic models (Cuaresma et al.,, 2004; Rahimiyan & Baringo,, 2016) are prevalent electricity price forecasting methods, but deep learning approaches like normalizing flows (e.g., Cramer et al., (2023)) are also gaining attention. Electricity prices forecasts enable DR computations in future applications, offering advantages over historical data, especially for long horizons and volatile markets (Nowotarski & Weron,, 2018). Unlike data-averaging techniques, price forecasting methods are not restricted to representing data of the present and recent past. Instead, they can incorporate trends of the future, although commonly restricted to the use of historical data and small forecasting windows (e.g., Alonso et al., (2016); Shah et al., (2021)). Notably, however, both historical data based as well as complex scenario forecasting methods have been rather inaccurate at predicting the long-term evolution of electricy prices (Zareipour et al.,, 2010; Weron,, 2014; Gabrielli et al.,, 2022), since this evolution is subject to policy changes and political events (Yang et al.,, 2017).
Publications on DSM of industrial processes, including our own studies, such as Caspari et al., (2019); Mucci et al., (2023); El Wajeh et al., (2024), heavily rely on the chosen electricity price profiles that must represent current trends and allow for generalizable conclusions. Selecting a single historical profile or averaging historical data favors simplicity. Clearly, averaged price trends do not fully capture all features of historical profiles, as price averaging yields a single and exclusive price scenario. Additionally, data averaging does not account for the temporal evolution of prices to consider future price trends. Nonetheless, simple and sufficiently representative price profiles are usually adequate for evaluation of the theoretical potential of DR applications. Conversely, sophisticated scenario generation methods provide multiple price scenarios, valuable for active market participation. Yet, these methods involve a more complex implementation, while forecasts of the long-term price trends are still subject to high uncertainty. Moreover, an extensive evaluation of all possible price scenarios, product demand situations, and other possible operating disturbances quickly results in a curse of dimensionality and is thus often beyond the scope of research studies on DR.
Herein, we aim to establish a simple method to construct meaningful and representative DA and ID benchmark price profiles to be used within future economic studies, e.g., cost optimal planning and scheduling, power-to-X, and energy-flexible design and operation. Within this method, we desire an option to ensure comparability between the DA and ID prices, which is useful for DR across multiple electricity markets, e.g., in bidding strategies (Silva et al.,, 2022), or scheduling and control (Caspari et al.,, 2020). This property is not provided by existing methods, e.g., Maciejowska et al., (2019); Teichgraeber & Brandt, (2019). The developed price profile will exhibit the following essential properties:
  1. 1.
    Method simple to understand, deterministic, and few parameters,
  2. 2.
    Evolution of electricity prices closely represents current price trends,
  3. 3.
    Comparability of DA and ID price profiles,
  4. 4.
    Magnitude of price fluctuations is adjustable (scalable), while 1. - 3. still hold.
To achieve the desired properties, we build our method on previous works on constructing average price profiles (Schäfer et al.,, 2020; Germscheid et al.,, 2022) and introduce adaptations tailored to the applications investigated. Because the focus of DR studies is on the economic potential of electricity price fluctuations rather than steady-state optimization, the price variance over the investigation period is assumed to hold greatest importance for optimal economic decision-making. Thus, our method allows for adjusting the magnitude of fluctuations in the price profile. Additionally, we ensure comparability between the DA and ID scenarios based on the near-zero mean distribution of the ID-DA deviation by equalizing the integrated DA and ID prices over the designated period. To underpin this approach, we provide statistical analyses of hourly DA and quarter-hourly real-time ID prices in historical data from European spot markets. Our approach can extend to other markets and price granularities (e.g., half-hourly) where similar observations hold.
In the remainder of this article, we first introduce a method for constructing a representative 24-hour DA price profile from historical data of the last available 365 days (Section 2.1). Next, we utilize the same historical data to extract the ID price profile, based on ID - DA price deviations (Section 2.2). Then, we introduce a few modifications and apply the same approach to generate single-week (Section 3) representative profiles. Last, we create ready-to-use benchmark DA and ID price profiles for future studies on DSM on the different time-scales presented (Section 4). This open-access data set is also provided via Git (Link).
2   Single-day price profile
2.1 DA price profile
We begin with the construction of the benchmark DA profile of one representative day. To this end, we select a price data set of 365 consecutive days to exclude seasonal effects (Knittel & Roberts,, 2005). The 365 days horizon may be chosen arbitrarily and if desired extended to multiple years. We remark that incidents and trends contained in the price data set, e.g., caused by political events and developments, will affect the price scenario created. Here, we use up-to-date price values of the full year 2023. Similar to our prior works (Schäfer et al.,, 2020; Germscheid et al.,, 2022), we first average the historical DA prices for every day hour, $k=1,2,...,24$, individually111 Notice that for $k\geq 2$, Eq. (1) may be equivalently rewritted as $\overline{DA}_{k}=\overline{DA}_{k-1}+\overline{\Delta DA}_{k}$, where $\overline{\Delta DA}_{k}=1/N\sum_{i=1}^{N}{(DA_{i,k}-DA_{i,k-1})}$. Hence, this profile captures the moving average price development over a day. :
$\overline{DA}_{k}=\frac{1}{N}\sum_{i=1}^{N}{DA_{i,k}}\,,$ (1)
where $N=365$ is the number of days of selected historical data, $DA_{i,k}$ is the DA price at hour $k$ of day $i=1,2,...,N$, and $\overline{DA}_{k}$ is the average DA price of hour $k$. If desired, days with significant outliers, weekends or holidays may be excluded. However, here we consider the full year.
Next, we calculate the overall mean value $\overline{DA}$:
$\overline{DA}=\frac{1}{24}\sum_{k=1}^{24}\overline{DA}_{k}\,,$
which is also the mean over all historical data.
Fig. 1: Average and scaled average ($\beta=1.47$) 24-hour DA price profile generated from historical 2023 data (EPEX,, 2023).
To facilitate the adjustment of the price fluctuation around the mean value (see Requirement 4), i.e., variance scaling, we introduce a scaling factor $\beta>0$. This aspect extends our prior works. We scale all $\overline{DA}_{k}$ around $\overline{DA}$ by $\beta$:
$\overline{DA}_{k}^{\beta}=\overline{DA}+\beta\cdot(\overline{DA}_{k}-\overline{DA})\,.$ (2)
Clearly, setting $\beta=1$ corresponds to no scaling. The resulting scenario exhibits low deviating behavior according to the historical data (Fig. 2). To minimize scenario complexity, we can thus use the $\beta=1$-profile as a moderate scenario. Importantly, the mean value $\overline{DA}$ remains identical irrespective of scaling.
As the daily standard deviation of the average profile might not represent the daily standard deviation of all data due to smoothing of the data (see Schäfer et al., (2020)), we use $\beta$ to correct the averaged results. Thus, in order to obtain a nominal benchmark profile, we choose $\beta$ such that:
$\sigma(\overline{DA}_{1}^{\beta},...,\overline{DA}_{24}^{\beta})=\frac{1}{N}\sum_{i=1}^{N}\sigma(DA_{i,1},...,DA_{i,24})\,,$
i.e., the standard deviation $\sigma$ of the resulting scaled price profile matches the average standard deviation of the daily DA price profiles. We insert the definition of the standard deviation:
$\sqrt{\frac{1}{24}\sum_{k=1}^{24}(\overline{DA}_{k}^{\beta}-\overline{DA})^{2}}=\frac{1}{N}\sum_{i=1}^{N}\sqrt{\frac{1}{24}\sum_{k=1}^{24}\big{(}DA_{i,k}-\frac{1}{24}\sum_{j=1}^{24}DA_{i,j}\big{)}^{2}}$
and obtain an explicit solution after mathematical reformulation:
$\beta=\frac{1}{N}\sum_{i=1}^{N}\sqrt{\frac{\sum_{k=1}^{24}\big{(}DA_{i,k}-\frac{1}{24}\sum_{j=1}^{24}DA_{i,j}\big{)}^{2}}{\sum_{k=1}^{24}(\overline{DA}_{k}-\overline{DA})^{2}}}\,.$
For example, for the full year 2023 and EPEX spot market, we receive the value $\beta=1.47$. The resulting DA profile is illustrated by Fig. 1. Here, the characteristic double peak profile is visible with minimal values occurring during the 4th and 14th hour as well as maximal values during the 8th and the 9th. The absolute price range after scaling ($\beta=1.47$) is $103.46\,\mathrm{€/MWh}$ with a mean value of $\overline{DA}=95.18\,\mathrm{€/MWh}$.
If a more distinct price variation is desired, we can specify a larger value of $\beta$. In order to systematically choose a larger $\beta$ factor that corresponds to some extreme behavior of the historical data (Requirement 2), we represent the density distribution of the standard deviations $\sigma(DA_{i,1},...,DA_{i,24})$ across all days $i=1,2,...,N$ through a histogram, Fig. 2. Subsequently, we identify a quantile indicative of high fluctuations within the historical data. We exemplary show the quantile 3.4 ($Q_{3.4}$), corresponding to the 15% highest values of the daily standard deviations, Fig. 2. The value of the quantile is the corresponding standard deviation, namely:
$\sigma(\overline{DA}_{1}^{\beta},...,\overline{DA}_{24}^{\beta})=Q_{3.4}\big{(}\{\sigma(DA_{i,1},...,DA_{i,24})\}_{i=1}^{N}\big{)}\,,$
satisfied for:
$\beta=\frac{Q_{3.4}\big{(}\{\sigma(DA_{i,1},...,DA_{i,24})\}_{i=1}^{N}\big{)}}{\sqrt{\frac{1}{24}\sum_{k=1}^{24}(\overline{DA}_{k}-\overline{DA})^{2}}}\,,$
giving $\beta=1.97$ for the full year 2023. Other options could include the choice of a preset high value, e.g., as an expression of the scaled nominal profile $\beta$, here $\beta=2.35$. Conversely, for a less aggressive profile, we may set $\beta<1$ to match, e.g., $Q_{0.8}$ (of the 20% least deviating historical data).
Fig. 2: Distribution histogram of daily DA price standard deviation. Based on historical data of year 2023.
2.2 ID price profile
We proceed with generating an average ID profile by first exploring the relationship between DA and ID prices, followed by the construction of an ID profile based on a DA reference. Rahimiyan & Baringo, (2016) assessed the correlation between DA and ID prices, pinpointing uncorrelated DA price and market deviation (also referred to as DA-ID price deviation or residual price). As noticed by Germscheid et al., (2022), the ID price oscillates around the DA price in a well-defined, season-independent harmonic pattern with dominant frequencies at $0.5\,\mathrm{h^{-1}}$ and close to $1\,\mathrm{h^{-1}}$. Consequently, representative ID profiles may be constructed by first subtracting ID and DA data and then using this market deviation to construct a harmonic deviation profile. Here, we build on these findings and introduce some further refinements. We consider the $\text{ID}_{\text{3}}$ price index, which is the volume-weighted average price of all trades that took place from three hours to 30 min before delivery (EPEX,, 2023).
We first calculate the average market deviation of ID around the DA profile for each 15 min interval of the historical data ($\overline{\Delta}_{q}^{ID}$) similar to Germscheid et al., (2022):
$\begin{split}\Delta_{i,q}^{ID}&=ID_{i,q}-DA_{i,\lceil q/4\rceil}\,,\\ \overline{\Delta}_{q}^{ID}&=\frac{1}{N}\sum_{i=1}^{N}\Delta_{i,q}^{ID}\,,\end{split}$ (3)
[…]
Fig. 3: Average ID-DA price deviation $\overline{\Delta}_{q}^{ID}$ for the full year 2023. Mean value indicated by dashed line.
Therein, we observe the deviation of an hourly pattern as indicated by a Fourier analysis (Germscheid et al.,, 2022). Notably, large jumps in the price deviation occur at full hours, typically involving a change of sign. Moreover, we observe phase shifts at hours 3, 7, 13, and 19, where the inner hour trend switches between continuously increasing and decreasing. Next, a straightforward approach for constructing an ID profile would be (Germscheid et al.,, 2022):
$\overline{ID}_{q}=\overline{DA}_{\lceil q/4\rceil}+\overline{\Delta}_{q}^{ID}$ (4)
However, to meet Requirement 3 we refine this approach by imposing the additional condition of zero cumulative market deviation:
$\sum_{q=1}^{96}\overline{\Delta}_{q}^{ID}=0\,,$ (5)
such that the cumulative prices of the benchmark DA and ID profiles will be identical. Figure 4(a) reveals that the daily cumulative deviation scaled by the daily cumulative DA price in 2023 is approximately symmetrically distributed around zero, which supports Eq. (5). Moreover, even the hourly integral of the ID to DA price deviations scaled by the hourly cumulative DA price appears to be symmetric around zero mean, Fig. 4(b). This observation is consistent with the dominant harmonic frequencies at $0.5\,\mathrm{h^{-1}}$ and $1\,\mathrm{h^{-1}}$ mentioned above. We found similar distributions for the years 2019 to 2022 (figures omitted for brevity). The distributions, although approximately zero-mean, exhibit significant deviations. Hence, they only valid to first order.
Fig. 4: Distribution histogram of relative cumulative daily (a) and hourly (b) deviation of ID and DA prices. Based on historical data of year 2023.
Because the distribution in Fig. 4(a) is, however, not perfectly Gaussian and exhibits non-zero mean (2.23 €/MWh, also see Fig. 3), Eq. (3) does not satisfy the comparability requirement (Eq. (5)). Hence, we modify the approach and extend Eq. (4) by an offset correction term to obtain:
$\tilde{\Delta}_{q}^{ID}=\overline{\Delta}_{q}^{ID}-\frac{1}{96}\sum_{q=1}^{96}\overline{\Delta}_{q}^{ID}\,.$ (6)
We notice that the correction approach may be further refined based on the zero-mean distribution in Fig. 4(b). However, we decide against such an additional refinement to keep the method simple. Next, similar to $\beta$ for the DA profile, we scale $\tilde{\Delta}^{ID}_{q}$ by a parameter $\gamma>0$ to match the year-average of the ID standard deviations:
$\sigma(\overline{ID}_{1}^{\gamma}...,\overline{ID}_{96}^{\gamma})=\frac{1}{N}\sum_{i=1}^{N}\sigma({ID}_{i,1},...,{ID}_{i,96})\,,$ (7)
wherein the benchmark ID profile is calculated as:
$\overline{ID}_{q}^{\gamma}=\overline{DA}_{\lceil q/4\rceil}^{\beta}+\gamma\tilde{\Delta}_{q}^{ID}\,,\,q=1,...,96\,.$
The nominal parameter value $\gamma$ is specified by solving the implicit nonlinear equation with $\beta$ fixed:
$\sqrt{\sum_{q=1}^{96}(\beta(\overline{DA}_{\lceil q/4\rceil}-\overline{DA})+\gamma\tilde{\Delta}_{q}^{ID})^{2}}=\frac{1}{N}\sum_{i=1}^{N}\sqrt{\sum_{q=1}^{96}\big{(}{ID}_{i,q}-\frac{1}{96}\sum_{q=1}^{96}{ID}_{i,q}\big{)}^{2}}\,.$ (8)
Equation (8) is readily solved numerically, e.g., using Microsoft Excel. For the full year 2023, we determine $\gamma=1.91$.
Again, we may adjust the fluctuation magnitude of an ID profile around its DA reference through a modification of $\gamma$. For example, when specifying $\gamma$ such that the standard deviation matches the quantile 3.4 of the historical ID standard deviation density distribution:
$\begin{split}Q_{3.4}\big{(}\{\sigma(ID_{i,1},...,ID_{i,96})\}_{i=1}^{N}\big{)}&=\sigma(\overline{ID}_{1}^{\gamma},...,\overline{ID}_{96}^{\gamma})\\ &=\sqrt{\frac{1}{96}\sum_{q=1}^{96}(\beta(\overline{DA}_{\lceil q/4\rceil}-\overline{DA})+\gamma\tilde{\Delta}_{q}^{ID})^{2}}\,,\end{split}$ (9)
we obtain $\gamma=1.77$ for 2023. We here note that although the $\gamma$-value in the extreme scenario is smaller than that of the nominal scenario, the overall deviation in the extreme ID profile is higher, due to the higher $\beta$-value of the extreme DA profile taking part in the calculations (Eq. (9)). Similar to the DA prices, larger and smaller values of $\gamma$ create more aggressive and less aggressive ID profile, respectively.
Figure 5 illustrates the final single-day nominal benchmark DA and ID profile based on full year 2023 data and for $\beta=1.47$ and $\gamma=1.91$. For the DA profile we note that prices exhibit a daily minimum, coinciding with peak photovoltaic generation around noon (Schäfer et al.,, 2020; Kiesel & Paraschiv,, 2017). Prices start rising around 6:00, when the workday begins and start falling around 18:00 when the workday ends. Two price peaks around 7:00 and 18:00 denote a high electricity demand meeting a low renewable energy supply from photovoltaics (Knittel & Roberts,, 2005). On the other hand, ID prices drop subhourly between 07:00 to 13:00 and 19:00 to 03:00 and rise in the meantime. Namely, while the sun is rising (from 07:00 to 13:00), offer is higher than the hourly average demand and thus price is lower in the last quarter of the hour. During afternoon hours (from 13:00 to 19:00), the offer is higher in the first quarter of the hour. Further, during offpeak hours (from 19:00 to 07:00) the oscillation of the average ID price follows the demand of power-intensive industry (decreasing steps), and the production design of the fossil power plants (increasing steps) for minimum production regulations (Kiesel & Paraschiv,, 2017).
Fig. 5: Nominal ($\beta=1.47$, $\gamma=1.91$) single-day DA and ID price profile for year 2023 (EPEX,, 2023).
3   Single-week price profile
3.1 DA price profile
Commonly, the time horizon of planning and scheduling spans more than one day of operation. In some cases, a periodic scenario is sufficient, wherein the price profile is duplicated for multiple consecutive days. However, if distinct prices are desired, the procedure from the previous section can be extended to price scenarios comprising multiple days. Here, we construct such a price profile for a one-week time window. The choice of this horizon length aligns with the significant DA price frequencies pinpointed by Germscheid et al., (2022), allowing us to capture and average over weekly patterns. However, we do not recommend to apply the averaging method to construct time frames longer than a week. In that case, the averaging would involve either statistically too small historical data sets (for $N=365$) or data sets spanning more than a year and thus including long-term changes of market dynamics and policy. However, one remedy to construct a profile spanning multiple weeks could be the periodic repetition of the averaged profile.
Following the approach presented in Section 2.1, we first average the historical DA prices for every hour $16824d+k=1,2,...,$ 168 of the week after splitting the available data in an integer number of full weeks $W$ (Monday to Sunday), so that $7\cdot W\leq N$:
$\overline{DA}_{24d+k}=\frac{1}{W}\sum_{w=1}^{W}{DA_{(d+1)+7(w-1),k}}\,,$
where $d=0,1,...,6$ indicates the day of the week $w=1,...,W$ (e.g, Monday corresponds to 0), $k=1,2,...,24$ denotes the hour of the day and $24d+k$ is the hour of the constructed single-week profile. Similar to Section 2.1 we use the hourly data of 2023, thus considering $W=52$.
We next calculate the overall mean value:
$\overline{DA}=\frac{1}{168}\sum_{k=1}^{24}\sum_{d=0}^{6}\overline{DA}_{24d+k}\,,$
and scale the profile as in Eq. (2) by factor $\beta$:
$\overline{DA}_{24d+k}^{\beta}=\overline{DA}+\beta\cdot(\overline{DA}_{24d+k}-\overline{DA})\,.$
For the construction of a nominal benchmark profile, $\beta$ is selected such that the standard deviation of the constructed single-week profile is equal to the average standard deviation of all weekly profiles of the historical data:
$\sigma(\overline{DA}_{1}^{\beta},...,\overline{DA}_{168}^{\beta})=\frac{1}{W}\sum_{w=1}^{W}\sigma(DA_{7w-6,1},...,DA_{7w-6,24},DA_{7w-5,1},...,DA_{7w,24})\,,$
which leads to the explicit $\beta$ expression:
$\beta=\frac{1}{W}\sum_{w=1}^{W}\sqrt{\frac{\sum_{d=0}^{6}\sum_{k=1}^{24}\big{(}DA_{(d+1)+7(w-1),k}-\frac{1}{168}\sum_{d=0}^{6}\sum_{k=1}^{24}DA_{(d+1)+7(w-1),k}\big{)}^{2}}{\sum_{d=0}^{6}\sum_{k=1}^{24}(\overline{DA}_{24d+k}-\overline{DA})^{2}}}\,.$ (10)
For more extreme price fluctuations, we select $\beta$ so that the standard deviation of the constructed profile is equal to the quantile 3.0 value of the density distribution:
$\beta=\frac{Q_{3.0}\big{(}\{\sigma(DA_{7w-6,1},...,DA_{7w-6,24},DA_{7w-5,1},...,DA_{7w,24})\}_{w=1}^{W}\big{)}}{\sqrt{\frac{1}{168}\sum_{d=0}^{6}\sum_{k=1}^{24}(\overline{DA}_{24d+k}-\overline{DA})^{2}}}\,.$ (11)
We here use a smaller quantile compared to the one selected in Section 2.1, because weeks with extremely varying prices do not usually follow the nominal fluctuating pattern, see also figures of the supplementary information (SI).
Considering the EPEX Spot DA market data for the year 2023, Eq. (10) gives $\beta=1.58$, and Eq. (11) $\beta=1.76$. More or less aggressive profiles are again obtained by adjusting the parameter $\beta$.
3.2 ID price profile
We extend our methodology on extracting a representative ID profile directly following the analysis in Section 2.2 to compose the ID profile over a one-week horizon by superposing the average ID-DA price deviation with the average DA price over a single-week period. Therefore, we first calculate the average market deviation of the historical data:
$\begin{split}\overline{\Delta}_{96d+q}^{ID}&=\frac{1}{W}\sum_{w=1}^{W}\big{(}ID_{(d+1)+7(w-1),q}-DA_{(d+1)+7(w-1),\lceil q/4\rceil}\big{)}\,,\\ \end{split}$
where $q=1,...,96$ is the quarter hour index, $ID_{(d+1)+7(w-1),q}$ is the ID price on week $w$, day of the week $d$ and quarter hour $q$.
As opposed to the hourly and daily historical profiles, the weekly profiles do not show a zero-mean distribution of the weekly cumulative ID-DA deviation. Thus, a zero-mean correction for the full week as in Eq. (6) is not expedient here. Instead, we apply Eq. (6) separately for each of the seven days of the week to enforce a daily closure:
$\sum_{q=1}^{96}\overline{\Delta}_{96d+q}^{ID}=0\,,\hskip 12.91663ptd=0,...,6$
and therefore:
$\tilde{\Delta}_{96d+q}^{ID}=\overline{\Delta}_{96d+q}^{ID}-\frac{1}{96}\sum_{q=1}^{96}\overline{\Delta}_{96d+q}^{ID}\,,\hskip 12.91663ptd=0,...,6\,.$
As a result, the single-week profile fulfills Requirement 3.
Again, the final ID profile is constructed by the superposition of the $\gamma$ scaled market deviations to the single-week constructed DA profile:
$\overline{ID}_{96d+q}^{\gamma}=\overline{DA}_{24d+\lceil q/4\rceil}^{\beta}+\gamma\tilde{\Delta}_{96d+q}^{ID}\,,\hskip 12.91663ptd=0,...,6\,,\,q=1,...,96\,.$
As in Eq. (7), we select $\gamma$ by solving the equation:
$\begin{split}\frac{1}{W}\sum_{w=1}^{W}\sigma(ID_{7w-6,1},...,ID_{7w-6,96},&ID_{7w-5,1},...,ID_{7w,96})\\ &=\sqrt{\sum_{d=0}^{6}\sum_{q=1}^{96}\big{(}\beta(\overline{DA}_{24d+\lceil q/4\rceil}-\overline{DA})+\gamma\tilde{\Delta}_{96d+q}^{ID}\big{)}^{2}}\end{split}$ (12)
for $\gamma>0$. In this way, the constructed single-week ID profile reproduces the standard deviation of the averaged standard deviation of the historical single-week ID profiles. For a stronger higher varying profile, the parameter $\gamma$ can be calculated as:
$\begin{split}Q_{3.0}\big{(}\{\sigma(ID_{7w-6,1},...,ID_{7w,96})\}_{w=1}^{W}\big{)}=\sqrt{\sum_{d=0}^{6}\sum_{q=1}^{96}\big{(}\beta(\overline{DA}_{24d+\lceil q/4\rceil}-\overline{DA})+\gamma\tilde{\Delta}_{96d+q}^{ID}\big{)}^{2}}\,.\end{split}$ (13)
Considering the historical data of EPEX Spot for 2023, and 52 full weeks (from Monday to Sunday), Eq. (12) gives $\gamma=1.60$, and Eq. (13) gives $\gamma=2.32$. The final single-week nominal benchmark DA and ID profiles for the full year 2023 data ($\beta=1.58$,$\gamma=1.60$) are shown in Fig. 6.
Figure 6 is consistent with the observations of Germscheid et al., (2023). Lowest prices are reached during the weekend associated with a lower demand. Accordingly, the price peak in the weekend mornings is less significant than the peak in the evening. As reported by Guthrie & Videbeck, (2002), during weekdays the daytime off-peak and evening peak periods exhibit distinct market behaviors. However, during the weekend these periods operate as a unified market. Additionally, the absence of morning peaks in the weekends results from morning prices correlated with the rest prices of the weekend. The highest DA prices are reached on Monday and Thursday evening, and the highest ID prices on Monday evening.
Fig. 6: Nominal ($\beta=1.58$, $\gamma=1.60$) single-week DA and ID price profile for year 2023.
4   Exemplary scenarios and comparison to historical data
We apply the proposed method to calculate DA and ID price profiles based on the full year 2023 historical data (01.01.2023 to 31.12.2023). We construct profiles following Sections 2 and 3 with three variance scaling cases each:
  1. 1.
    Single-day horizon:
    i) Nominal ($\beta=1.47$, $\gamma=1.91$), ii) Moderate ($\beta=1.0$, $\gamma=1.0$), iii) Extreme ($\beta=1.85$, $\gamma=1.77$),
  2. 2.
    Single-week horizon:
    i) Nominal ($\beta=1.58$, $\gamma=1.6$), ii) Moderate ($\beta=1.0$, $\gamma=1.0$), iii) Extreme ($\beta=1.76$, $\gamma=2.32$).
We show the resulting single-day nominal profiles in Fig. 5 and provide the moderate and extreme profile in the SI. Similarly, the single-week nominal profiles are shown in Fig. 6 and the moderate and extreme profiles can be found in the SI. We furthermore provide the profiles in CSV format via Git (Link) as well as in table format in the SI. We highlight that the Moderate data may be used as a basis to construct price profiles for arbitrary scaling options ($\beta$, $\gamma$).
(a) DA profiles
(b) ID profiles
Fig. 7: Nominal single-day and corresponding best fit of historical profiles for year 2023.
(a) DA profile
(b) ID profile
Fig. 8: Moderate single-day and corresponding best fit of historical profile for year 2023.
(a) DA market
(b) ID market
Fig. 9: Nominal single-week price profiles compared to the best fit historical week in 2023.
For the sake of comparison, we complement the constructed profiles by historical profiles that closely match each profile. The historical profiles are retrieved by minimizing the mean absolute deviation between the historical and a constructed scenario for DA and ID simultaneously. In the single-day case, we obtain:
$\min_{i}\Big{(}\sum_{k=1}^{24}|\overline{DA}_{k}^{\beta}-DA_{i,k}|+\frac{1}{4}\sum_{q=1}^{96}|\overline{ID}_{q}^{\gamma}-ID_{i,q}|\,.\Big{)}$ (14)
A single-week historical profile is identified analogously. Figure 7 contrasts the nominal single-day profiles and the historical day with the best fit. While the inner hour variance of the nominal ID profile is more pronounced during early and late hours, all profiles follow similar trends and are overall comparable. Moreover, inserting Eq. (2) into Eq. (14) reveals that the historic profile in 7 indeed reflects the day which deviates the least from the rest historical days, i.e., the cluster center in this sense. Next, Fig. 8 depicts the moderate profiles and the historical best fit. Since here all single-week profiles have the same historical week as the best fit, we only show the plot for the nominal profile in Fig. 9. This approach is equivalent to determining the cluster center of a single cluster formed by all available time series. We conclude that the moderate and nominal profile represent the historical data closely. For the nominal profile, the fitted historical day merely serves for validation, whereas for the moderate profile the historical day corresponds to the most representative day in the historical data set.
Next, we discuss a limitation of our approach. Compared to the close alignment of the moderate and nominal profiles with historical data, solving the fitting problem in Eq. (14) for the extreme profile did not lead to satisfactory results. This is explained by the extreme deviations or outliers observed in the historical data, especially in the ID price. These outliers, although leading to a high variance, which is considered for when composing the extreme profiles, exhibit different patterns than the averaged ones of the constructed (scaled) profiles. Nevertheless, when fitting historical DA and ID profiles separately rather than jointly, we were able to find similar historical profiles (see figures in SI). We conclude that while a more extreme variance scaling is generally valid, an alternative scenario generation method should be consulted if extreme profiles with abnormal patterns are sought.
In addition to the qualitative assessment of the profiles by the figures, we provide a quantitative statistical evaluation in Tables 4 to 4. In particular, the tables confirm that the standard deviation of the constructed nominal DA and ID profiles and the corresponding historical data match as intended by nominal scaling. Moreover, while the mean of the DA profiles matches the historical data, there is a difference for the ID profiles due to the correction of the integral zero-mean deviation. Consequently, the 24 h price integrals of the corresponding DA and ID profiles are identical (compare last row in Tables 4 and 4 as well as Tables 4 and 4).
Table 1: Characteristic statistical values of the single-day DA electricity price profile.
Average value over all week integrals
🟊 The number of histogram bins is calculated with the Scott rule (Scott,, 1979)
Average value over all full weeks
Quantity DA price profile
Nominal Moderate Extreme historical
$\beta$ 1.47 1.00 1.85🟊 -
Minimum (€/MWh) 54.99 67.82 67.82 $-500.00$
Maximum (€/MWh) 158.45 138.24 138.24 524.27
Mean (€/MWh) 95.18 95.18 95.18 95.18
Std. deviation (€/MWh) 28.22 19.21 35.46 28.22
24 h integral (€/24MWh) 2284.21 2284.21 2284.21 2284.21
Table 2: Characteristic statistical values of the single-day ID electricity price profile.
Quantity ID price profile
Nominal Moderate Extreme historical
$\gamma$ 1.91 1.00 1.77 -
Minimum (€/MWh) 16.36 49.04 12.00 -666.87
Maximum (€/MWh) 202.96 161.50 215.73 3543.51
Mean (€/MWh) 95.18 95.18 95.18 97.40
Std. deviation (€/MWh) 40.31 24.61 44.73 40.31
24 h integral (€/24MWh) 2284.21 2284.21 2284.21 2337.64
Table 3: Characteristic statistical values of the single-week DA electricity price profile.
Quantity DA price profile
Nominal Moderate Extreme historical
$\beta$ 1.58 1.00 1.76🟊 -
Minimum (€/MWh) -22.07 20.92 -35.70 -500.00
Maximum (€/MWh) 191.34 156.22 202.47 524.27
Mean (€/MWh) 95.40 95.40 95.40 95.40
Std. deviation (€/MWh) 38.59 24.46 43.07 38.59
1 week integral (€/24MWh) 16026.52 16026.52 16026.52 16026.52
Table 4: Characteristic statistical values of the single-week ID electricity price profile.
Quantity ID price profile
Nominal Moderate Extreme historical
$\gamma$ 1.60 1.00 2.32🟊 -
Minimum (€/MWh) -64.06 -5.18 -101.93 -666.87
Maximum (€/MWh) 279.91 211.45 336.05 3543.51
Mean (€/MWh) 95.40 95.40 95.40 95.61
Std. deviation (€/MWh) 47.02 29.69 57.48 47.02
1 week integral (€/24MWh) 16026.52 16026.52 16026.52 16026.52
5   Conclusions
We present a method to construct electricity price scenarios based on an averaging and scaling procedure using historical price data. Our approach is capable of generating both DA and ID scenarios and the cumulative deviation of a [DA, ID] pair can be adjusted to zero-mean deviation, if comparability is needed. Our method is constructed to be simple and intuitive, to have few adjustable parameters, and thus be easily reproducible, while at the same time providing sufficiently representative price profiles for future case studies on DSM.
We have applied our method to historical DA and ID data from EPEX spot for the year 2023. There, we have generated multiple price scenarios with varying magnitude of price fluctuation. The comparison to historical data confirmed that our proposed method is capable of providing representative price profiles of adjustable variance for both DA and ID markets simultaneously. Moreover, the constructed profiles automatically satisfy a set of basic Requirements 1 to 4. Consequently, the tedious and non-deterministic procedure of selecting a representative historical profiles while ensuring the desired profile properties hold is circumvented.
Clearly, some advanced scenario generation methods also consider higher-order statistical moments in their generation procedure. However, these approaches are inherently more involved and result in a greater number of adjustable parameters. In general, both historical data based as well as complex scenario forecasting methods have been rather inaccurate at predicting the long-term evolution of prices, as this evolution is subject to policy and political changes. Hence, we focus on approaches to reflect the current character of the electricity price on spot markets. Since we are interested in an intuitive and deterministic method rather than covering all possible price profile shapes, we accept to trade such a flexibility for the desired simplicity of the proposed averaging-based method.
Acknowledgements. The authors gratefully acknowledge the financial support of the Kopernikus project SynErgie by the Federal Ministry of Education and Research (BMBF), and the project supervision by the project management organization Projektträger Jülich (PtJ). The authors gratefully acknowledge the financial support of and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)—333849990/GRK2379 (IRTG Hierarchical and Hybrid Approaches in Modern Inverse Problems). Moreover, we thank Sonja Germscheid and Eike Cramer for constructive feedback on the manuscript.
Supplementary information
Chrysanthi Papadimitriou, Jan C Schulze, and Alexander Mitsos
 Abstract. Herein, we provide supplementary material to the listed publication.
Extreme & Moderate DA & ID profiles
(a) Moderate
(b) Extreme
Fig. 1: Moderate ($\beta=1.00$, $\gamma=1.00$) and Extreme ($\beta=1.85$, $\gamma=1.77$) single-day DA and ID price profile for year 2023.
(a) Moderate
(b) Extreme
Fig. 2: Moderate ($\beta=1.00$, $\gamma=1.00$) and Extreme ($\beta=1.76$, $\gamma=2.32$) single-week DA and ID price profile for year 2023.
Best fits of historical single-day and single-week profiles
(a) DA market
(b) ID market
Fig. 3: Extreme single-day price profiles compared to the best fit historical day in 2023.
(a) DA market
(b) ID market
Fig. 4: Extreme single-week price profiles compared to the best fit historical week in 2023.
Fig. 5: Extreme DA single-week price profiles compared to one of the best fitted historical weeks in 2023.
Data of single-day DA and ID profiles
Table 1: Profiles of hourly DA electricity prices.
Hour DA price in €/MWh Hour DA price in €/MWh
$k$ Nominal Moderate Extreme $k$ Nominal Moderate Extreme
1 78.81 84.03 74.61 13 58.92 70.50 49.63
2 72.84 79.97 67.11 14 54.99 67.82 44.69
3 69.47 77.68 62.89 15 63.81 73.83 55.77
4 68.52 77.04 61.69 16 80.34 85.07 76.53
5 74.05 80.80 68.64 17 103.81 101.05 106.02
6 93.27 93.88 92.78 18 133.73 121.42 143.61
7 115.06 108.71 120.15 19 158.45 138.24 174.67
8 120.63 112.5 127.15 20 150.24 132.65 164.35
9 112.18 106.75 116.54 21 129.49 118.53 138.29
10 94.45 94.68 94.27 22 113.27 107.49 117.91
11 80.35 85.08 76.55 23 99.93 98.41 101.15
12 69.27 77.54 62.63 24 88.33 90.51 86.57
Table 2: Profiles of 15 min ID electricity prices.
Quarter ID price in €/MWh Quarter ID price in €/MWh
$q$ Nominal Moderate Extreme $q$ Nominal Moderate Extreme
1 105.11 97.78 98.88 49 83.31 83.24 72.13
2 86.65 88.13 81.84 50 59.63 70.87 50.28
3 69.46 79.15 65.99 51 48.22 64.91 39.75
4 52.32 70.20 50.18 52 28.78 54.75 21.82
5 90.66 89.28 83.56 53 35.71 57.75 26.91
6 73.83 80.49 68.03 54 45.66 62.95 36.08
7 65.65 76.22 60.48 55 53.25 66.92 43.09
8 57.75 72.09 53.19 56 65.72 73.43 54.59
9 75.31 80.73 68.27 57 16.36 49.04 12.00
10 65.99 75.86 59.67 58 52.29 67.81 45.15
11 63.81 74.72 57.66 59 77.41 80.93 68.32
12 62.63 74.10 56.57 60 103.50 94.56 92.39
13 60.35 72.76 54.15 61 20.45 53.79 21.29
14 62.29 73.78 55.94 62 69.7 79.52 66.72
15 69.40 77.49 62.50 63 99.36 95.02 94.09
16 76.46 81.18 69.02 64 129.15 110.58 121.57
17 55.09 70.89 51.14 65 42.19 68.86 49.18
18 64.02 75.55 59.38 66 89.87 93.77 93.17
19 81.40 84.64 75.42 67 125.09 112.17 125.65
20 95.86 92.19 88.75 68 162.54 131.73 160.20
21 58.88 75.91 61.06 69 79.43 93.05 93.52
22 81.30 87.62 81.74 70 115.21 111.74 126.53
23 103.84 99.40 102.53 71 149.51 129.66 158.17
24 121.73 108.75 119.04 72 198.27 155.13 203.15
25 93.37 97.38 100.14 73 143.59 130.48 160.96
26 112.29 107.26 117.60 74 143.09 130.22 160.50
27 130.14 116.58 134.06 75 167.54 142.99 183.05
28 120.20 111.40 124.90 76 202.96 161.50 215.73
29 156.87 131.43 160.58 77 199.54 158.41 209.83
30 118.76 111.52 125.43 78 150.58 132.83 164.66
31 111.31 107.63 118.55 79 139.01 126.79 153.99
32 92.53 97.82 101.23 80 131.18 122.7 146.77
33 149.08 126.02 150.58 81 187.53 148.85 191.83
34 116.79 109.16 120.79 82 136.91 122.41 145.13
35 95.71 98.15 101.35 83 119.83 113.49 129.38
36 69.39 84.39 77.06 84 104.97 105.72 115.66
37 139.50 118.22 135.82 85 158.76 131.26 159.88
38 105.77 100.60 104.71 86 121.80 111.95 125.78
39 80.90 87.60 81.76 87 104.83 103.08 110.12
40 49.30 71.09 52.61 88 78.47 89.31 85.80
41 117.6 104.54 110.91 89 137.52 118.05 135.82
42 87.21 88.67 82.87 90 109.54 103.43 110.01
43 70.87 80.13 67.81 91 93.70 95.16 95.40
44 42.27 65.19 41.42 92 67.40 81.42 71.14
45 100.50 93.86 91.45 93 120.31 107.22 116.08
46 72.70 79.34 65.80 94 99.44 96.32 96.82
47 61.04 73.24 55.03 95 78.06 85.15 77.10
48 38.55 61.5 34.29 96 57.24 74.27 57.89